# Surfel-based Gaussian Inverse Rendering for Fast and Relightable Dynamic Human Reconstruction from Monocular Videos

### v2

@@ 4.4.2 Comparison on real-world dataset
We also compare our method with the baselines on the real-world nearly A-pose dataset qualitatively in Fig. 5. The novel poses are selected from the AIST dataset [107]. We can observe that the albedo estimation of R4D* [25] fails to recover the plausible solution, its geometry is inaccurate. The relighting results at both the training and test poses are unsuitable for the environmental light. Compared with Intrinsic Avatar [29], our methods can produce more plausible relighting results in addition to the 5×5$\times$5 × faster in training speed and 100×100$\times$100 × faster in rendering speed.
@@ 4.4.3 PBR properties estimation results on the real-world datasets with complex motions
We follow the setting in the InstantNVR [34] and choose view 4 from ZJU-MoCap as the training view. We visualize the intrinsic estimation results in Fig. 6. Our methods can produce plausible BRDF estimation results. We also render the clothed human at novel illumination and get well-reproduced results. We also visualize the approximated occlusion calculated from the template mesh to show the effectiveness of our proposed human avatar occlusion approximation strategy. We can observe from Fig. 6 that our method can achieve a plausible PBR properties estimation from a monocular video with complex human motions, and can also re-render them with an appearance suitable to the environment lighting.
@@ 4.5.1 Training strategy
  Method Albedo Normal Error $\downarrow$
@@
We first compare different training strategies for the PBR optimization stage in the Tab. II and the Fig. 8. ”Freeze shape” means that we choose the optimization strategy of GS-IR [47], which freeze the shape after image reconstruction stage and only optimize the PBR materials of each Gaussian. ”w/o progressive” means that we do not utilize the strategy in (equation 21) and directly utilize the normal-consistency loss (equation 19) from the first stage. We can observe from the Fig. 8 that if we utilize the ”freeze shape”, the geometry of Gaussian is fixed. The geometry from the first stage cannot bring any improvement. If we do not utilize our proposed progressive training strategy, the normal of each Gaussian is always enforced to align with the inaccurate geometry in the first stage. The inaccurate geometry will hurt the geometry optimization in the PBR stage. With our proposed progressive training strategy, both the normal of each splat and the geometry of the avatar can be aligned consistently to the accurate geometry. Moreover, the inaccurate geometry also leads to inaccurate albedo estimation. The quantitative results in Tab. II further demonstrate these.
4.5.2 Primitive representation
@@
  Method Albedo Normal Error $\downarrow$
@@
We ablate different choices of primitive representation in Tab. III and Fig. 9. Since original 3DGS [32] do not have an attribute for splat orientation, we follow the previous relightable 3DGS [47, 50] to assign a learnable normal vector for each Gaussian and keep all the other strategy unchanged. We can observe from Fig. 9 that with 3DGS as a primitive representation, though the splat normal (R-normal) can be very smooth, there exists a lot of uneven region in the Depth (D-normal). The R-normal and the D-Normal are not consistent. However, with the 2DGS representation, the R-Normal and the D-Normal can be both consistent and accurate. The Normal Error in Tab. III also demonstrates this quantitatively.
@@ 4.5.3 Normal-consistency loss
  Method Albedo Normal Error $\downarrow$ Relighting (novel pose)
@@
We ablate the normal consistency loss in Tab IV and Fig. 10. ”$\ell_{1}$-loss” means that we follow the previous relightable 3DGS [50, 47] and utilize the $\ell_{1}$-loss to minimize the difference of the linear blended normal and the normal from depth points. We keep the remaining factors unchanged. We can observe that utilizing the $\ell_{1}$-loss, the normal of each splat cannot align with the D-Normal well, while the inaccurate splat normal also leads to bad relighting results as shown in the Fig. 10. The quantitative results from Tab IV also provide evidence.
@@
  Method Albedo Normal Error $\downarrow$
@@ 4.5.5 Occlusion calculation from template mesh or clothed mesh
We ablate our approximation of the occlusion calculation from the template mesh in Tab. VI and Fig. 12. For the fully clothed mesh case, we follow 2DGS [33] and use the TSDF Fusion [109] to extract the mesh of the clothed human avatar at the canonical pose. We query the points by using tri-linear interpolation from the canonical skinning weight voxel to get the skinning weight for each canonical vertex. We keep other regularization terms the same as those in 2DGS and optimize the same steps.Tab. VI shows that the accuracy of the estimated PBR properties with the occlusion calculated from the template mesh only has a minor gap to the occlusion calculated from the fully clothed mesh. However, the rendering time cost using the fully clothed mesh is about 10×10$\times$10 × longer than that with the template mesh. This is because the number of vertices and faces in the fully clothed mesh can be 20×20$\times$20 × more than those in the template mesh. Fig. 12 shows that the calculated visibility map and the estimated albedo do not have any significant difference. This demonstrates that our proposed occlusion approximation strategy can achieve a faster rendering speed without sacrificing the accuracy of the estimated PBR properties.
@@
  Method Albedo Rendering Time $\downarrow$
@@ 4.5.6 PBR regularization
We ablate our PBR regularization choices in the Tab. VII and Fig. 13. ”w/o. white reg” means that we do not utilize the white light regularization in Equation 23. ”w/o. smooth reg” means that we do not utilize the smooth regularization in Equation 22. We can observe from Fig. 13 that without the light regularization, the albedo estimation and the splat normal can be noisy. Without the smooth regularization, the albedo estimation does not include details (such as eyes on the face), the splat normal and the roughness are totally messy. The roughness value varied even in the region that should have the same material properties. The PBR rendering image may also fit concisely with the ground-truth image. The quantitative results in Tab. VII also demonstrate the effectiveness of regularization terms.
@@
  Method Albedo Rendering Time $\downarrow$
@@
5 Conclusion And Future Work
@@
While our method efficiently reconstructs the PBR properties of clothed human avatars from monocular videos, it is important to acknowledge its limitations. One significant limitation is the lack of detailed facial expression capture. Our approach does not leverage expression capture techniques, resulting in less detailed facial animations for the avatars. Future work could enhance realism and expressiveness by integrating facial expression techniques, as explored in studies such as [110, 111, 112], with our full-body gesture reconstruction. A promising direction is to combine the expressive Gaussian Avatar representation with our full PBR representation pipeline to achieve both relightable and expressive human avatar reconstruction.

### v3

@@ 4.4.2 Comparison on real-world dataset
We also qualitatively compare our method with the baselines on the dataset of nearly A-pose in the real world in Fig. 5. The novel poses are selected from the AIST dataset [112]. We can observe that the albedo estimation of R4D* [25] fails to recover the plausible solution, its geometry is inaccurate. The relighting results at both the training and test poses are unsuitable for environmental light. Compared with Intrinsic Avatar [29], our methods can produce more plausible relighting results in addition to $5\times$ faster in training speed and $100\times$ faster in rendering speed.
4.4.3 Comparision with non-PBR 3DGS-based baseline
We compared our method with GART  [39], a non-PBR baseline for human reconstruction in the Fig. 8. Although our method consumes more training time, it can successfully reconstruct all the PBR properties of the clothed avatar and support the relighting under novel illuminations. Moreover, as demonstrated in the Tab. VI, the occlusion calculation is the main bottleneck for time consumption compared with the non-PBR baseline. Our training speed improvements benefit from our proposed occlusion calculation.
Figure 8: Comparisions with Non-PBR baseline GART [39] . Our method achieves PBR physical properties including albedo, normal, metallics and roughness, while GART can only support radiance fields reconstruction.
@@ 4.4.4 PBR properties estimation results on the real-world datasets with complex motions
We follow the setting in the InstantNVR [34] and choose view 4 from ZJU-MoCap as the training view. We visualize the intrinsic estimation results in Fig. 6. Our methods can produce plausible BRDF estimation results. We also render the clothed human at novel illumination and get well-reproduced results. We also visualize the approximated occlusion calculated from the template mesh to show the effectiveness of our proposed human avatar occlusion approximation strategy. We can observe from Fig. 6 that our method can achieve a plausible PBR properties estimation from a monocular video with complex human motions, and can also re-render them with an appearance suitable to the environment lighting. Moreover, since our proposed occlusion calculation strategy is based on the SMPL human model, it cannot calculate the occlusion correctly at the complex geometric region, therefore leading to the wrong physical properties at the winkle area as highlighted by the red rectangles in Fig. 6. We also provide a comprehensive discussion about this at the Section  6.
@@ 4.4.5 Reposed clothed human avatars in novel views under novel illuminations
  Method Albedo
Normal
Error
$\downarrow$
@@
4.5 Ablation studies
4.5.1 Training strategy
@@
  Method Albedo
Normal
Error
$\downarrow$
@@
We first compare different training strategies for the PBR optimization stage in the Tab. II and the Fig. 9. ”Freeze shape” means that we choose the optimization strategy of GS-IR [46], which freeze the shape after image reconstruction stage and only optimize the PBR materials of each Gaussian. ”w/o progressive” means that we do not utilize the strategy in (equation 21) and directly utilize the normal-consistency loss (equation 19) from the first stage. We can observe from the Fig. 9 that if we utilize the ”freeze shape”, the geometry of Gaussian is fixed. The initial geometry from the first stage does not improve anything. Without our progressive training, the Gaussian normals would always align with the inaccurate geometry from the first stage, hindering geometry optimization in the PBR stage. With our proposed progressive training strategy, both the normal of each splat and the geometry of the avatar can be aligned consistently to the accurate geometry. Moreover, the inaccurate geometry also leads to inaccurate albedo estimation. The quantitative results in Tab. II further demonstrate these.
4.5.2 Primitive representation
We ablate different choices of primitive representation in Tab. III and Fig. 10. Since original 3DGS [32] do not have an attribute for splat orientation, we follow the previous relightable 3DGS [46, 49] to assign a learnable normal vector for each Gaussian and keep all the other strategies unchanged. We can observe from Fig. 10 that with 3DGS as a primitive representation, though the splat normal (R-normal) can be very smooth, there exists a lot of uneven region in the Depth (D-normal). The R-normal and the D-Normal are not consistent. However, with the 2DGS representation, the R-Normal and the D-Normal can be both consistent and accurate. The Normal Error in Tab. III also demonstrates this quantitatively.
@@ 4.5.3 Normal-consistency loss
  Method Albedo
Normal
Error
$\downarrow$
Relighting (novel pose)
@@
We ablate the normal consistency loss in Tab IV and Fig. 11. $\ell_{1}$-loss” means that we follow the previous relightable 3DGS [49, 46] and utilize the $\ell_{1}$-loss to minimize the difference of the linear blended normal and the normal from depth points. When we keep the other factors unchanged and use the $\ell_{1}$-loss, the splat normals cannot align well with the D-Normals. This inaccurate splat normal also leads to poor relighting results, as shown in Fig. 11. The quantitative results in the Tab IV further support this observation.
@@
  Method Albedo
Normal
Error
$\downarrow$
@@ 4.5.5 Occlusion calculation from template mesh or clothed mesh
[…]
@@
  Method Albedo
Rendering
Time
$\downarrow$
@@ 4.5.6 PBR regularization
We ablate our PBR regularization choices in the Tab. VII and Fig. 14. ”w/o. white reg” means that we do not utilize the white light regularization in Equation 23. ”w/o. smooth reg” means that we do not utilize the smooth regularization in Equation 22. From Fig. 14, we can see that without light regularization, the albedo and splat normals are noisy. Without smoothness regularization, the albedo lacks details, and the splat normals and roughness are chaotic, with inconsistent roughness even in regions of the same material. This leads to PBR renders that don’t closely match the ground truth. The quantitative results in Tab VII also show the effectiveness of these regularization terms.
@@
  Method Albedo
Rendering
Time
$\downarrow$
@@
5 Limitation And Future Work
While our method efficiently reconstructs the PBR properties of clothed human avatars from monocular videos, it is important to acknowledge its limitations.
1). One significant limitation is the lack of detailed facial expression capture. Our approach does not leverage expression capture techniques, resulting in less detailed facial animations for the avatars. Future work could enhance realism and expressiveness by integrating facial expression techniques with our full-body gesture reconstruction, as explored in studies such as [115, 116, 117]. A promising direction is to combine the expressive Gaussian Avatar representation with our full PBR representation pipeline to achieve both relightable and expressive human avatar reconstruction.
2). Another severe limitation comes from the SMPL model. Although the SMPL model effectively achieves a highly efficient occlusion approximation for successfully estimating physical properties, it cannot model wrinkles, loose clothes, or other detailed geometric deformations caused by clothes since it only models the unclothed human mesh. Exploring other detailed human models [118, 119] for the occlusion approximation is important to achieve a detailed occlusion calculation.
3). Lastly, since our method focuses on the reconstruction of the PBR properties from monocular videos, our current PBR-based inverse rendering can only provide one possible solution as most previous works [29, 25]. Scale ambiguity may exist between the reconstructed PBR properties and the unknown illumination light intensity. Incorporating the diffusion-based intrinsic estimation method [120, 121, 122] is a promising research direction to solve the scale ambiguity issue.
6 Conclusion